# Electromagnetic field in a cavity induced by gravitational waves

### v1

@@ III.3 Perturbation of Elasticity
Considering static material at the leading-order, we set $v^{a}=u^{a}$ and $\chi_{ab}=\gamma_{ab}$. Perturbations of the normalization condition Eq. 8, the Lorentz factor Eq. 21, and the spatial velocity Eq. 22 provide the following:
@@
Eqs. 23 and 13 for the material metric give
@@
We introduce the perturbed stress-energy of the material given by
@@
We choose $\tilde{T}\left(\epsilon\right)=O\left(\epsilon^{2}\right)$ to be compatible with the perturbed Einstein equation in Eq. 60 and assume that there is no other contribution except the material on $\tilde{T}\left(\epsilon\right)$ up to $\epsilon^{3}$. Considering homogeneous material without stress at the leading-order, we set $\nabla_{a}\rho=0$, and $\sigma_{ab}=0$. Then, Eq. 16 becomes
@@
Note that $\delta\sigma$ is spatial and gauge-invariant because its leading-order vanishes.
@@
where $D\cdot X$ is the divergence for a spatial vector $X$. Differentiating Eq. 86 by $\nabla_{u}$, substituting Eq. 78, and utilizing Eqs. 70 and 71, we obtain
@@
The Helmholtz decomposition allows the separation of solenoidal and irrotational modes for $\delta V$. The solenoidal mode (S wave), satisfying $D\cdot\left(\delta V\right)_{\mathrm{S}}=0$, and the irrotational mode (P wave), satisfying $D\times\left(\delta V\right)_{\mathrm{P}}=0$, have wave equations, respectively, as follows:
@@ III.4 Perturbation of EM Field
Because we imposed $\tilde{F}\left(\epsilon\right)=O\left(\epsilon^{2}\right)$, the EM contribution on $\tilde{T}\left(\epsilon\right)$ is at $O\left(\epsilon^{4}\right)$. It is compatible with the perturbed Einstein’s equations Eq. 60 and the perturbed contracted Bianchi identities Eq. 86. For the leading-orders, we set $j_{a}=0$ and $F_{ab}=\tensor{\varepsilon}{{}^{c}_{ab}}B_{c}$ where $B$ is the constant magnetic field. Notice that $\delta j$ is gauge-invariant because its leading-order vanishes. Then, the linear perturbation of Eqs. 27, 28 and 29 becomes:
@@
Note that $\delta E$ is spatial and gauge-invariant because its leading-order vanishes. By contracting $u$ with Eqs. 92 and 94, we obtain equations for $\delta E$ as
@@ III.5 Inside the Cavity
To derive the perturbations of conditions Eqs. 32 and 33 within the cavity, let us introduce $\mathcal{W}_{\epsilon}$ as the 4-dimensional volume for the cavity motion in $\mathcal{M}_{\epsilon}$. By considering a gauge $\phi_{\epsilon}$ such that $\phi_{\epsilon}\left[\mathcal{W}_{0}\right]=\mathcal{W}_{\epsilon}$, the perturbations of Eqs. 32 and 33 on $\mathcal{W}_{0}$ can be expressed as follows:
@@
Because $\delta V$ and $\delta j$ are gauge-invariant, these conditions hold in any gauge. Subsequently, within the cavity, Eqs. 96 and 97 transform into the following equations:
@@ III.6 Gauges and Rigidity
Eqs. 100, 101 and 102 hold for all gauges. Moreover, each term in these equations is gauge-invariant as they involve only gauge-invariant quantities. The contribution of GWs comes from only the right-hand side in Eq. 101. This term vanishes when the direction of GWs aligns with the magnetic field direction due to Eq. 64. This observation contradicts the claims made in Berlin et al. (2022); Domcke et al. (2022). Let us discuss the reason.
In Berlin et al. (2022), they impose rigidity of the material by fixing the 4-velocity for conductor elements $v$ as $\left(\partial\middle/\partial t\right)$ in the proper detector frame even when GWs pass. Then, $\delta v=0$ due to Eq. 53 in the proper detector gauge, and Eqs. 92 and 94 become
@@
In these equations, the second term on the right-hand side in Eq. 103 represents the “effective charge”, while the third, fourth, and fifth terms on the right-hand side in Eq. 104 constitute the temporal derivative of spatial “effective current”, as defined in Berlin et al. (2022).
We would like to address three points. Firstly, $\left(\partial/\partial t\right)$ in the perturbed spacetime $\mathcal{M}_{\epsilon}$ is not normalized. To accurately obtain the electric field, it is essential to use a normalized vector in Eq. 30. Secondly, $\delta v$ is a gauge-dependent quantity due to $\mathcal{L}_{\xi}v\neq 0$ in $\mathcal{M}_{0}$. Consequently, $\delta v$ is not controllable experimentally. Enforcing $\delta v=0$ requires a condition imposed by Eq. 76, such as $\left(\delta V\right)^{a}=-\left(\gamma\delta u\right)^{a}$, where $\delta u$ is the gauge-dependent perturbation in the proper detector gauge. This cannot be the solution for Eq. 87 that is the acoustic wave equation with its propagating velocity smaller than the speed of light, while $\delta u$ has the phase of GWs. Thirdly, a perfect rigid body does not exist in the framework of general relativity because it violates causality, as discussed in Giulini (2010).
@@ IV.1 Electric Field inside Cavity
where we include the dissipation term with the quality factor $Q_{n}$ and $f_{n}\left(t\right)$ is the external “force” defined by
@@
where $\mathfrak{e}^{a}_{n}\left(\vec{x}\right)$ is the metric dual to $\mathfrak{e}^{n}_{a}\left(\vec{x}\right)$. Notice that the factor $\mathcal{V}/\omega_{n}^{2}$ in Eq. 106 makes the dimension of “force” identical to the dimension of energy over the electric field.
@@
×$\times\int d^{2}\kappa\,\tilde{h}_{ab}\left(\omega,\kappa\right)\left(\hat{B}\times\kappa\right)^{a}\bar{\mathfrak{e}}_{n}^{b}\left(\omega,\kappa\right)e^{i\omega\kappa\cdot\vec{x}_{0}}.$ (109)
Here, ~$\tilde{\mathcal{E}}_{n}\left(\omega\right)$ and $\tilde{f}_{n}\left(\omega\right)$ are the Fourier transformations of $\mathcal{E}_{n}\left(t\right)$ and $f_{n}\left(t\right)$, respectively. $A\left(\omega;\omega_{n},Q_{n}\right)$ and $a\left(\omega;\omega_{n},Q_{n}\right)$ are defined in Eqs. 42 and 43. Also, $\left|B\right|\equiv\sqrt{B\cdot B}$, $\hat{B}\equiv B/\left|B\right|$, and $\times$ is the cross product with $\varepsilon$. Additionally, $\vec{x}_{0}$ is a position inside $\mathcal{V}$, and $\bar{\mathfrak{e}}_{n}\left(\omega,\kappa\right)$ is defined by
@@ IV.2 Signal and Antenna Pattern
In the context of GW detection using mode $n$, we introduce the GW signal $h\left(t\right)$, the detector tensor $\tilde{D}^{ab}\left(\omega,\kappa\right)$ and the transfer function $\tilde{T}\left(\omega\right)$ as follows:
@@
We present the expression for the GW amplitude following the format in Section 7 of Park (2022):
@@
Here, $H\left(\omega,\kappa\right)>0$ represents the amplitude strength, $\chi\left(\omega,\kappa\right)$ is the ellipticity, $\psi\left(\omega,\kappa\right)$ is the polarization angle, $\xi\left(\omega,\kappa\right)$ is the phase. The real orthonormal basis $\left\{e^{+},e^{\times}\right\}$ is defined by:
@@
where $\Lambda_{abcd}\equiv\sum_{A}e^{A}_{ab}e^{A}_{cd}$. Note that this $\mathfrak{F}\left(\omega,\lambda\right)$ does not depend on the ellipticity $\chi$.
IV.3 Example: Cylindrical Cavity with TM010 Mode
@@
For a cylindrical cavity, we find that $\bar{\mathfrak{e}}_{n}^{a}\left(\omega,\kappa\right)$ defined in Eq. 110 is real or pure imaginary because of the parity symmetry, as in Eq. 38. Therefore, the antenna pattern given in Eq. 126 is applicable to the cylindrical cavity. We introduce a pair of polar and azimuthal angles $\left(\theta,\phi\right)$ for the GW propagation such that $\kappa\left(\theta,\phi\right)=\sin\theta\cos\phi\hat{x}+\sin\theta\sin\phi\hat{y}+\cos\theta\hat{z}$ and parameter $\alpha$ for the magnetic field direction as $\hat{B}\left(\alpha\right)=\cos\alpha\hat{z}+\sin\alpha\hat{x}$. Then, $\bar{\mathfrak{e}}_{010}\left(\omega,\kappa\right)$ for TM010 mode using Eq. 110 is given by
@@
$\bar{\mathfrak{e}}_{010}\left(\omega,\theta,\phi\right)=2\frac{\sin\left(\frac{1}{2}\omega L\cos\theta\right)}{\frac{1}{2}\omega L\cos\theta}\frac{\mathfrak{j}_{0,1}\mathfrak{J}_{0}\left(\omega R\sin\theta\right)}{\mathfrak{j}^{2}_{0,1}-\left(\omega R\sin\theta\right)^{2}}\hat{z},$ (127)
where the origin of $\vec{x}^{\prime}$ in Eq. 110, i.e. $\vec{x}_{0}$ in Eq. 109, is selected as the center of the cylinder, and $L$ denotes the length of the cylinder. From Eq. 126, we get the antenna pattern as
@@
$\mathfrak{F}\left(\omega,\theta,\phi\right)$ $=\frac{1}{2}\left|\bar{\mathfrak{e}}_{010}\left(\omega,\theta,\phi\right)\right|$
@@
×$\times\left[\sin^{4}\theta-\cos\theta\sin^{3}\theta\cos\phi\sin\left(2\alpha\right)\right.$
$\quad\left.+\sin^{2}\theta\left\{1+\sin^{2}\theta\left(\sin^{2}\phi-2\right)\right\}\sin^{2}\alpha\right]^{\frac{1}{2}}.$ (128)
Fig. 1 shows antenna patterns in the resonance frequency for several $\alpha$ and $L/R=1$. If the direction of GWs is parallel to the cylinder axis or the magnetic field, there is no GW signal, as shown in the Fig. 1.
Figure 1: This figure depicts the antenna patterns, denoted as $\mathfrak{F}\left(\omega_{010},\theta,\phi\right)$ in Eq. 128, corresponding to magnetic field directions $\alpha\in\left\{0,\pi/6,\pi/3,\pi/2\right\}$ and an aspect ratio $L/R=1$. The gray arrows indicate the magnetic field directions. Antenna patterns for $\alpha>\pi/2$ can be obtained by rotating the patterns in the figure by $\pi$ around the axis of cylinder and reversing the direction of the gray arrows. This symmetry arises from the relation $\mathfrak{F}\left(\omega,\theta,\phi;\alpha\right)=\mathfrak{F}\left(\omega,\theta,\pi+\phi;\pi-\alpha\right)$.

### v2

@@ III.3 Perturbation of elasticity
Considering static material at the leading order, we set $v^{a}=u^{a}$ and $\chi_{ab}=\gamma_{ab}$. Perturbations of the normalization condition Eq. 8, the Lorentz factor Eq. 21, and the spatial velocity Eq. 22 provide the following:
@@
Equations 23 and 13 for the material metric give
@@
We introduce the perturbed stress–energy tensor of the material given by
@@
We choose $\tilde{T}\left(\epsilon\right)=O\left(\epsilon^{2}\right)$ to be compatible with the perturbed Einstein equation in Eq. 72 and assume that there is no other contribution except the material on $\tilde{T}\left(\epsilon\right)$ up to $\epsilon^{3}$. Considering homogeneous material without stress at the leading order, we set $\nabla_{a}\rho=0$, and $\sigma_{ab}=0$. Then, Eq. 16 becomes
@@
Note that $\delta\sigma$ is spatial and gauge invariant because its leading order vanishes.
@@
where $D\cdot X$ is the divergence for a spatial vector $X$. Differentiating Eq. 98 by $\nabla_{u}$, substituting Eq. 90, and utilizing Equations 82 and 83, we obtain
@@
The Helmholtz decomposition allows the separation of solenoidal and irrotational modes for $\delta V$. The solenoidal mode (S wave), satisfying $D\cdot\left(\delta V\right)_{\mathrm{S}}=0$, and the irrotational mode (P wave), satisfying $D\times\left(\delta V\right)_{\mathrm{P}}=0$, where $D\times$ is the curl, have wave equations, respectively, as follows:
@@ III.4 Perturbation of EM field
Because we imposed $\tilde{F}\left(\epsilon\right)=O\left(\epsilon^{2}\right)$, the EM contribution on $\tilde{T}\left(\epsilon\right)$ is at $O\left(\epsilon^{4}\right)$. It is compatible with the perturbed Einstein’s equations Eq. 72 and the perturbed contracted Bianchi identities Eq. 98. For the leading orders, we set $j_{a}=0$ and $F_{ab}=\tensor{\varepsilon}{{}^{c}_{ab}}B_{c}$ where $B$ is the constant magnetic field. Notice that $\delta j$ is gauge invariant because its leading order vanishes. Then, the linear perturbation of Equations 27, 28 and 29 becomes:
@@
Note that $\delta E$ is spatial and gauge invariant because its leading order vanishes. By contracting $u$ with Equations 104 and 106, we obtain equations for $\delta E$ as
@@ III.5 Inside the cavity
To derive the perturbations of conditions Equations 32 and 33 within the cavity, let us introduce $\mathcal{W}_{\epsilon}$ as the four-dimensional volume for the cavity motion in $\mathcal{M}_{\epsilon}$. By considering a gauge $\phi_{\epsilon}$ such that $\phi_{\epsilon}\left[\mathcal{W}_{0}\right]=\mathcal{W}_{\epsilon}$, the perturbations of Equations 32 and 33 on $\mathcal{W}_{0}$ can be expressed as follows:
@@
Because $\delta V$ and $\delta j$ are gauge invariant, these conditions hold in any gauge. Subsequently, within the cavity, Equations 108 and 109 transform into the following equations:
@@ III.6 Gauges and rigidity
Equations 112, 113 and 114 hold for all gauges. Moreover, each term in these equations is gauge invariant as they involve only gauge-invariant quantities. The contribution of GWs comes from only the right-hand side in Eq. 113. This term vanishes when the direction of GWs aligns with the magnetic field direction due to Eq. 76. This observation contradicts the claims made in Berlin et al. (2022); Domcke et al. (2022). Let us discuss the reason.
In Berlin et al. (2022), they impose rigidity of the material by fixing the four-velocity for conductor elements $v$ as $\left(\partial\middle/\partial t\right)$ in the proper detector frame even when GWs pass. Then, $\delta v=0$ due to Eq. 65 in the proper detector gauge, and Equations 104 and 106 become
@@
In these equations, the second term on the right-hand side in Eq. 115 represents the “effective charge,” while the third, fourth, and fifth terms on the right-hand side in Eq. 116 constitute the temporal derivative of spatial “effective current,” as defined in Berlin et al. (2022).
We would like to address three points. Firstly, $\left(\partial/\partial t\right)$ in the perturbed spacetime $\mathcal{M}_{\epsilon}$ is not normalized. To accurately obtain the electric field, it is essential to use a normalized vector in Eq. 30. Secondly, $\delta v$ is a gauge-dependent quantity due to $\mathcal{L}_{\xi}v\neq 0$ in $\mathcal{M}_{0}$. Consequently, $\delta v$ is not controllable experimentally. Enforcing $\delta v=0$ requires a condition imposed by Eq. 88, such as $\left(\delta V\right)^{a}=-\left(\gamma\delta u\right)^{a}$, where $\delta u$ is the gauge-dependent perturbation in the proper detector gauge. This cannot be the solution for Eq. 99, which is the acoustic wave equation with its propagating velocity smaller than the speed of light, while $\delta u$ has the phase of GWs. Thirdly, a perfect rigid body does not exist in the framework of general relativity because it violates causality, as discussed in Giulini (2010).
@@ IV.1 Electric field inside cavity
where we include the dissipation term with the quality factor $Q_{n}$ following Berlin et al. (2022) and $f_{n}\left(t\right)$ is the external “force” defined by
@@
where $\mathfrak{e}^{a}_{n}\left(\vec{x}\right)$ is the metric dual to $\mathfrak{e}^{n}_{a}\left(\vec{x}\right)$. Notice that the factor $\mathcal{V}/\omega_{n}^{2}$ in Eq. 118 makes the dimension of “force” identical to the dimension of the energy over the electric field.
@@
$\qquad\int d^{2}\kappa\,\tilde{h}_{ab}\left(\omega,\kappa\right)\left(\hat{B}\times\kappa\right)^{a}\bar{\mathfrak{e}}_{n}^{b}\left(\omega,\kappa\right)e^{i\omega\kappa\cdot\vec{x}_{0}}.$ (121)
Here, ~$\tilde{\mathcal{E}}_{n}\left(\omega\right)$ and $\tilde{f}_{n}\left(\omega\right)$ are the Fourier transformations of $\mathcal{E}_{n}\left(t\right)$ and $f_{n}\left(t\right)$, respectively. In Eq. 120, $A\left(\omega;\omega_{n},Q_{n}\right)$ and $a\left(\omega;\omega_{n},Q_{n}\right)$ are defined in Equations 54 and 55. In Eq. 121, $\tilde{h}_{ab}$ is defined by
$\tilde{h}_{ab}\left(\omega,\kappa\right)\equiv 2\omega^{-2}\tilde{R}_{acbd}\left(\omega,\kappa\right)u^{c}u^{d},$ (122)
satisfying
$\tensor{\tilde{h}}{{}^{a}_{a}}\left(\omega,\kappa\right)$ $=0,$ (123)
$\tilde{h}_{ab}\left(\omega,\kappa\right)u^{b}$ $=0,$ (124)
$\tilde{h}_{ab}\left(\omega,\kappa\right)\kappa^{b}$ $=0.$ (125)
Note that $\tilde{h}_{ab}$ is identical to the amplitude of metric perturbation in transverse-traceless gauge; however, we consider it as a derived quantity from the gauge-invarant perturbation of the Riemann tensor as in Eq. (27.24) in Thorne and Blandford (2017). Also, $\left|B\right|\equiv\sqrt{B\cdot B}$, $\hat{B}\equiv B/\left|B\right|$, and $\times$ is the cross product with $\varepsilon$. Additionally, $\vec{x}_{0}$ is a position inside $\mathcal{V}$, and $\bar{\mathfrak{e}}_{n}\left(\omega,\kappa\right)$ is defined by
@@ IV.2 Signal and antenna pattern
In the context of GW detection using mode $n$, we introduce the GW signal $h\left(t\right)$, the detector tensor $\tilde{D}^{ab}\left(\omega,\kappa\right)$, and the transfer function $\tilde{T}\left(\omega\right)$ as follows:
@@
We present the expression for $\tilde{h}_{ab}$ following the format in Sec. 7 of Park (2022):
@@
Here, $H\left(\omega,\kappa\right)>0$ represents the amplitude strength, $\chi\left(\omega,\kappa\right)$ is the ellipticity, $\psi\left(\omega,\kappa\right)$ is the polarization angle, and $\xi\left(\omega,\kappa\right)$ is the phase. The real orthonormal basis $\left\{e^{+},e^{\times}\right\}$ is defined by:
@@
where $\Lambda_{abcd}\equiv\sum_{A}e^{A}_{ab}e^{A}_{cd}$. Note that this $\mathfrak{F}\left(\omega,\kappa\right)$ does not depend on the ellipticity $\chi$.
IV.3 Cylindrical cavity
@@
For a cylindrical cavity, we observe that $\bar{\mathfrak{e}}_{n}^{a}\left(\omega,\kappa\right)$ defined in Eq. 126 has the form of a real vector multiplied by a complex scalar because of the parity symmetry, as in Eq. 51. Therefore, the antenna pattern given in Eq. 136 is applicable to the cylindrical cavity. We introduce a pair of polar and azimuthal angles $\left(\theta,\phi\right)$ for the GW propagation such that $\kappa\left(\theta,\phi\right)=\sin\theta\cos\phi\hat{x}+\sin\theta\sin\phi\hat{y}+\cos\theta\hat{z}$ and parameter $\alpha$ for the magnetic field direction as $\hat{B}\left(\alpha\right)=\cos\alpha\hat{z}+\sin\alpha\hat{x}$. Then, we get the antenna pattern from Eq. 136 as
@@
$\mathfrak{F}^{\mathrm{TM}}_{mnps}\left(\omega,\vartheta,\varphi\right)=4\mathfrak{j}_{mn}\mathfrak{A}^{\mathrm{TM}}_{mnp}\left[1+\left(\frac{R}{L}\frac{p\pi}{\mathfrak{j}_{mn}}\right)^{2}\right]$
@@
×$\qquad\times\left|L\omega\frac{\mathfrak{C}_{\left(-1\right)^{p+1}}\left(L\omega\cos\vartheta/2\right)}{\left(L\omega\cos\vartheta\right)^{2}-\left(p\pi\right)^{2}}\right|$
×$\qquad\times\left|\frac{\mathfrak{J}_{m}\left(R\omega\sin\vartheta\right)}{\left(R\omega\sin\vartheta\right)^{2}-\mathfrak{j}_{mn}^{2}}\cos\left(\vartheta\right)\sin\left(\vartheta\right)\mathfrak{C}_{m,-s}\left(\varphi\right)\right|$
@@
×$\qquad\times\sqrt{1-\left(\cos\varphi\sin\vartheta\sin\alpha+\cos\vartheta\cos\alpha\right)^{2}},$ (137)
$\mathfrak{F}^{\mathrm{TE}}_{mnps}\left(\omega,\vartheta,\varphi\right)=4\mathfrak{j}^{\prime}_{mn}\mathfrak{A}^{\mathrm{TE}}_{mnp}\left|p\pi\frac{\mathfrak{C}_{\left(-1\right)^{p+1}}\left(L\omega\cos\vartheta/2\right)}{\left(L\omega\cos\vartheta\right)^{2}-\left(p\pi\right)^{2}}\right|$
×$\qquad\times\left[\left(\frac{m}{\mathfrak{j}^{\prime 2}_{mn}}\frac{\mathfrak{J}_{m}\left(R\omega\sin\vartheta\right)}{R\omega\sin\vartheta}\cos\vartheta\mathfrak{C}_{ms}\left(\varphi\right)\right)^{2}\right.$
$\qquad\qquad\left.+\left(\frac{\mathfrak{J}^{\prime}_{m}\left(R\omega\sin\vartheta\right)}{\left(R\omega\sin\vartheta\right)^{2}-\mathfrak{j}^{\prime 2}_{mn}}\mathfrak{C}_{m,-s}\left(\varphi\right)\right)^{2}\right]^{\frac{1}{2}}$
×$\qquad\times\sqrt{1-\left(\cos\varphi\sin\vartheta\sin\alpha+\cos\vartheta\cos\alpha\right)^{2}},$ (138)
where $\vartheta\equiv\pi-\theta$ and $\varphi\equiv\pi+\phi$.
Figures 1 and 2 show antenna patterns using $\mathrm{TM}_{010}$ and $\mathrm{TE}_{212-}$ modes, respectively, at the resonance frequency for $\alpha\in\left\{0,\pi/6,\pi/3,\pi/2\right\}$ and $L/R=1$. Antenna patterns for $\alpha>\pi/2$ can be obtained by rotating the patterns in the figures by $\pi$ around the axis of cylinder and reversing the direction of the gray arrows. This symmetry arises from the relation $\mathfrak{F}\left(\omega,\vartheta,\varphi;\alpha\right)=\mathfrak{F}\left(\omega,\vartheta,\pi+\varphi;\pi-\alpha\right)$. Figures 3 and 4 show cross sections of the antenna patterns on a plane that contains the magnetic field vectors, for $\mathrm{TM}_{010}$ and $\mathrm{TE}_{212-}$ modes, respectively. The panel (a) in Fig. 4 distinctly illustrates a difference between our antenna patterns and those presented in Berlin et al. (2022). Our results reflect the fact that there is no GW signal when the direction of GWs is parallel to the cylinder axis or the magnetic field.
Figure 1: This figure depicts the antenna patterns of cylindrical cavity with $\mathrm{TM}_{010}$ mode, given in Eq. 137, at its resonance frequency. The patterns correspond to magnetic field directions $\alpha\in\left\{0,\pi/6,\pi/3,\pi/2\right\}$ and an aspect ratio $L/R=1$. The gray arrows indicate the magnetic field directions. (a) $\alpha=0$. (b) $\alpha=\pi/6$. (c) $\alpha=\pi/3$. (d) $\alpha=\pi/2$.
Figure 2: This figure depicts the antenna patterns of cylindrical cavity with $\mathrm{TE}_{212-}$ mode, given in Eq. 138, at its resonance frequency. The patterns correspond to magnetic field directions $\alpha\in\left\{0,\pi/6,\pi/3,\pi/2\right\}$ and an aspect ratio $L/R=1$. The gray arrows indicate the magnetic field directions. (a) $\alpha=0$. (b) $\alpha=\pi/6$. (c) $\alpha=\pi/3$. (d) $\alpha=\pi/2$.
Figure 3: The figure illustrates the antenna patterns of cylindrical cavity with $\mathrm{TM}_{010}$ mode, given in Eq. 137, at its resonance frequency. The patterns are depicted on the plane of $\varphi=0$ and $\varphi=\pi$, corresponding to magnetic field directions $\alpha\in\left\{0,\pi/6,\pi/3,\pi/2\right\}$, with an aspect ratio $L/R=1$. The gray arrows indicate the magnetic field directions. (a) $\alpha=0$. (b) $\alpha=\pi/6$. (c) $\alpha=\pi/3$. (d) $\alpha=\pi/2$.
Figure 4: The figure illustrates the antenna patterns of cylindrical cavity with $\mathrm{TE}_{212-}$ mode, given in Eq. 138, at its resonance frequency. The patterns are depicted on the plane of $\varphi=0$ and $\varphi=\pi$, corresponding to magnetic field directions $\alpha\in\left\{0,\pi/6,\pi/3,\pi/2\right\}$, with an aspect ratio $L/R=1$. The gray arrows indicate the magnetic field directions. (a) $\alpha=0$. (b) $\alpha=\pi/6$. (c) $\alpha=\pi/3$. (d) $\alpha=\pi/2$.